Solve: Log 3/4 log 8 (x 2 +7) +log 1/2 log 1/4 (x 2 +7) -1 =-2

Solve:        
Log 3/4log 8 (x2+7) +log1/2 log1/4(x2 +7)-1 =-2

Grade:12

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
Please confirm if the question is:
\left [ log_{3/4}log_{8} \right \left ( x^{2}+7 \right )]+\left [ log_{1/2}log_{1/4} \right \left ( x^{2}+7 \right )^{-1}]=-2
Then you can simplifiy this as follows:
\left [ log_{3/4}\left ( \frac{log\left ( x^{2}+7 \right )}{3log2} \right ) \right ]+\left [ log_{1/2}\left ( \frac{log\left ( x^{2}+7 \right )}{2log2} \right ) \right ]=-2
If we assume:
u=\frac{log\left ( x^{2}+7 \right )}{log2}
Then the bove expression can be solved easily.
Regards
Sumit